You are here

Beam Delivery News

Primary tabs

TItle Date Excerpt Attached images
Cyclotron down due to water pump failure 13 August 2014

The RF system tripped off at about 06:15 this morning. On investigation, it was found that the north Cu LCW pump had failed, roughly halving the pressure in the LCW system. Ops has shut down other loads on the system, including the main magnet power supply, as a precaution against damage.

Beam delivery update 11 August 2014

At this moment, the cyclotron is running and delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4.

Beam delivery update: weekend edition 08 August 2014

The cyclotron is currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This is scheduled to continue through the weekend.

Beam delivery update 07 August 2014

The cyclotron is still delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4 for isotope production at STF; this is still scheduled to continue through the weekend, as is 11Be delivery to TIGRESS for experiment S1429.

Beam delivery update 06 August 2014

The cyclotron is running normally, delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. Operation is scheduled to continue like this through the weekend.

Beam delivery update: Operation restored 05 August 2014

Operation was restored at about 17:45 and the cyclotron is currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~97 µA to BL2C4. This is expected to continue until roughly 08:00 tomorrow, at which point BL2C will be switched to PIF until sometime in the afternoon.

Beam delivery update: Maintenance today 05 August 2014

It was decided, despite the late start, to go ahead with a source filament and backplate change today. This is expected to take ~6 hours. Beam went off just before 13:00 and is expected back early this evening.

Beam delivery update: weekend edition 04 July 2014

The cyclotron is under vacuum, the repairs to the STF beamline window have been completed, and the inflector and deflector have been successfully tested at their operating voltages with the main magnet off. At this point, and assuming no unforeseen problems between now and then, it looks like beam will be injected some time tomorrow (Saturday).

Beam delivery…er, cyclotron repair update 03 July 2014

The removable section of the vertical injection beamline has been replaced and the cyclotron is being baked out overnight; the plan is to start cooling it back down at 05:00 so that the Cryogenics group can begin cooling the tank cryopanels later in the morning.

Beam delivery update: lid down; cyclotron pumping 02 July 2014

The cyclotron tank lid has been lowered and pumping is getting underway. At this point, it’s estimated that we might inject Saturday for delivery on Sunday.

Pages

Subscribe to Group News